FTMW OBSERVATION AND ANALYSIS OF THE p-H_{2}–AgCl AND o-H_{2}–AgCl COMPLEX

\underline{G. S. GRUBBS II}, DANIEL A. OBENCHAIN, HERBERT M. PICKETT and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, 52 Lawn Avenue, Middletown, CT, 06459-0180, USA (email to GSG2: ggrubbs@wesleyan.edu).

The rotational spectrum of p-H_{2}–AgCl and o-H_{2}–AgCl has been measured for the first time using a Balle-Flygare type Fourier transform microwave (FTMW) spectrometer. \frac{(B+C)}{2} ’s, nuclear quadrupole coupling constants, and centrifugal distortion constants have been determined for multiple isotopologues of both species while spin-spin coupling constants have also been determined for at least one isotopologue of the o-H_{2} species. Substantial changes in the eQq value from the monomer^{a} occur at the Cl nucleus upon complexation with the H_{2} and will be discussed. Experimental r_{0}’s for the H_{2} C.O.M. distance to Ag and Ag distance to Cl are 1.809(2)Å and 2.2656(2)Å, respectively, for the p-H_{2} species and will be compared to theory. Quantum chemical calculations were performed with an APFD^{b} density functional and MP2 with an aug-cc-pVQZ basis set for the hydrogen and chlorine with the effective core potential ECP28MDF AVQZ^{c}^{,}^{d} for the Ag and will be presented.

aK. D. Hensel, C. Styger, W. Jäger, A. J. Merer, and M. C. L. Gerry, J. Chem. Phys., 99(1993) 3320.

bA. Austin, G. A. Petersson, M. J. Frisch, F. J. Dobek, G. Scalmani, and K. J. Throsselll. Chem. Theor. Comp., 8(2012) 4989.

cD. Figgen, G. Rauhut, M. Dolg, and H. Stoll. Chem. Phys., 311(2005) 227.

dK. A. Peterson and C. Puzzarini. Theor. Chem. Acc., 114(2005) 283.